# Hankun Wang1, Chenpeng Du1, Yiwei Guo1, Shuai Wang2, Xie Chen1, Kai Yu1

### v1

@@ Abstract
keywords:
Decoder-only Text-to-Speech, Self-Attention Mechanisms, Training-Free Optimization
@@ 1 Introduction
With the widespread use of discrete audio tokens [8, 9, 10], the research paradigm in language models [11, 12] has shown a profound impact on speech modeling and synthesis [13, 14]. Motivated by recent advancements in auto-regressive (AR) models employing decoder-only architectures for text generation [12, 15], several studies, such as VALL-E [3] and BASE TTS [16] apply similar architectures to TTS task. These studies demonstrate the remarkable capacity of decoder-only architectures in producing natural-sounding speech. Unfortunately, robustness is a challenging issue for such models, due to the lack of explicit monotonic alignment constraints and the gap between teacher-forced training and AR inference [17]. To mitigate this, VALL-T [18] proposes a decoder-only transducer for imposing monotonic alignment constraints. However, this method requires training a new model under a different criterion. Drawing from experience with prior TTS systems, we believe that existing decoder-only TTS models should also implicitly learn some form of alignment via the internal attention black box during training. If we can identify the internal representation of alignment within the model and apply forced alignment constraints during inference, then building a lightweight method to enhance model robustness becomes feasible. A key insight is that in the neural TTS model Tacotron [1, 19] with an encoder-decoder architecture, researchers discovered diagonal speech-text alignments within its sole attention module’s attention map. This suggests that alignments in decoder-only models might also manifest within their attention maps. However, in decoder-only TTS models that employ massive attentions across various layers and heads [3, 20], different attention maps are responsible for different functionalities and are not always diagonal. This complicates both awareness and control of the alignment from the attention maps.
@@
(c) AEAM in a repeat generation.
(d) AEAM in a skipping generation.
Figure 1: Illustration of real attention alignments. (a), (c) and (d) are attention heatmaps from the same AEAM of the same model, with different inputs. (a) is for a text-speech pair directly picked from the test set. With the same inputs, (b) shows the reference alignment produced by the Viterbi algorithm. (c) and (d) illustrate the AEAM when some text tokens are repeated or skipped in synthesis.
In our initial experiments on the popular decoder-only TTS model VALL-E, we observed from the attention maps that certain attention heads within a specific layer exhibit diagonal patterns, indicating the alignment between speech and text. The attention maps here specifically represent the attention from speech to text, where speech tokens function as queries while text tokens act as keys. We call the attention maps of those heads Alignment-Emerged Attention Maps (AEAMs, cf. Fig. 1(a)). In the circumstances of repeating and skip generation, the alignments on AEAMs also exhibit corresponding patterns (cf. Fig. 1(c) and 1(d)). Based on this discovery, we introduce a novel inference method named Attention-Constrained Inference (ACI), which facilitates monotonic synthesis without changing the model structure and the training process. It first identifies all AEAMs among the attention maps of all layers and heads using the Attention Sweeping algorithm and then applies constraining masks (CMasks) on AEAMs to guide the monotonic generation process. The main contributions of our work are listed below:
@@ 2.2 Attention Sweeping
$\mathcal{C}_{\mathrm{E}}(\mathbf{M})=-\frac{1}{T}\sum_{t=1}^{T}\sum_{\ell=1}^{L}M_{t,\ell}\cdot\log{M_{t,\ell}}.\vspace{-0.4em}$ (1)
@@
Figure 2: An example of AEAM attention values and DP values, and corresponding CMasks. Here $\bm{x}$ is the input text tokens, and $\bm{y}$ is the speech tokens. (a) presents an example of the attention value pattern of a normalized AEAM when generating $y_{t+1}$, with the maximum value of each row highlighted. The values are normalized so that each row sums up to $1$. The constraining mask (CMask) for the newly generated token $y_{t}$ is now needed. (b) shows the CMask obtained via the trivial argmax strategy, which is prone to causing instability, such as the attention center being falling back. Based on attention values in (a), subfigure (c) illustrates the DP values $d$ of each state. (d) is the CMask obtained via the DP strategy, which shows stability in maintaining the monotonicity.
@@
$d_{t,\ell}=\begin{cases}d_{t-1,1}+(m_{t}-1)^{2}&\ell=1,\\ \min\{d_{t-1,\ell},d_{t-1,\ell-1}\}+(m_{t}-\ell)^{2}&l\geq 2.\end{cases}\vspace{-0.6em}$ (2)
@@
$\mathcal{C}_{\mathrm{A}}(\mathbf{M})=\frac{1}{T}\left(\mathcal{E}(\bm{m},\bm{a},T)+\min_{c}\mathcal{E}(\bm{a}+c,\bm{b},T)\right)\vspace{-0.5em}$ (3)
@@
$\mathcal{C}_{\mathrm{E}}\left(\mathbf{M}\right)+\mathcal{C}_{\mathrm{A}}\left(\mathbf{M}\right)<2\tau.\vspace{-0.3em}$ (4)
@@ 2.3 Attention Constraining
VALL-E’s AR synthesis often encounters issues with word skips, repeats, or pronunciation errors. We notice that when the text and synthesized speech do not perfectly match, the attention alignment in AEAMs show correspoding patterns (cf. Fig. 1(c) and 1(d)). Assuming that $y_{1}$ to $y_{t}$ have been input or synthesized and that $y_{t+1}$ is to be generated, a possible AEAM value pattern is shown in Fig. 2(a). Note that the key in the last row is the newly generated $y_{t}$, for which the alignment position needs to be calculated in this AR iteration. However, the AEAM may mistakenly align $y_{t}$ with a phoneme identical to the true one but positioned differently, leading to a non-monotonic attention alignment that misguide the subsequent generation.
@@ 3 Experiment
Figure 3: Calculate $\mathcal{C}_{\mathrm{E}}$ and $\mathcal{C}_{\mathrm{A}}$ for each attention map of 4 models, and plot them as 2-D points. The separation line is set at $\tau=1$.
@@ 3.4 Ablation Study
Attn. Constrained AEAM(s) WER(%) $\downarrow$ Del(%) Ins(%) Sub(%)
(2-16) and (2-4) 4.14 0.65 0.65 2.84
(2-16) 4.28 0.68 0.71 2.89
all 2-nd layer attention maps 4.76 0.89 0.80 3.06
@@ 4 Conclusions
Our paper introduced a training-free method, ACI, to improve the decoder-only TTS model VALL-E’s synthesis robustness. Utilizing Attention Sweeping, we identify and filter AEAMs, strongly related to the alignment, then employ attention constraining to mitigate synthesis errors. Experimental results demonstrate noticeable improvements in WER without degrading naturalness. Our work demonstrates that decoder-only TTS models can unsupervisedly learn alignment-like structures and shows a way to locate and exploit them, laying the foundation for various potential works, such as streaming generation or integrate additional signals into attention maps during training.

### v2

# Attention-Constrained Inference for Robust Decoder-Only Text-to-Speech

@@ Abstract
Index Terms— Decoder-only Text-to-Speech, Self-Attention Mechanisms, Training-Free Optimization
@@ 1 Introduction
With the widespread use of discrete audio tokens [8, 9, 10], the research paradigm in language models [11, 12] has shown a profound impact on speech modeling and synthesis [13, 14]. Motivated by recent advancements in auto-regressive (AR) models employing decoder-only architectures for text generation [12, 15], several studies, such as VALL-E [3] and BASE TTS [16] apply similar architectures to TTS task. These studies demonstrate the remarkable capacity of decoder-only architectures in producing natural-sounding speech. Unfortunately, robustness is a challenging issue for such models, due to the lack of explicit monotonic alignment constraints and the gap between teacher-forced training and AR inference [17]. To mitigate this, VALL-T [18] proposes a decoder-only transducer for imposing monotonic alignment constraints. However, this method requires training a new model under a different criterion. Drawing from experience with prior TTS systems, we believe that existing decoder-only TTS models should also implicitly learn some form of alignment via the internal attention black box during training. If we can identify the internal representation of alignment within the model and apply forced alignment constraints during inference, then building a lightweight method to enhance model robustness becomes feasible. A key insight is that in the neural TTS model Tacotron [1, 19] with an encoder-decoder architecture, researchers discovered diagonal speech-text alignments within its sole attention module's attention map. This suggests that alignments in decoder-only models might also manifest within their attention maps. However, in decoder-only TTS models that employ massive attentions across various layers and heads [3, 20], different attention maps are responsible for different functionalities and are not always diagonal 1. This complicates both awareness and control of the alignment from the attention maps.
(a) Head 1 of layer 1
(b) Head 4 of layer 2
(c) Head 3 of layer 3
(d) Head 3 of layer 4
(e) Head 1 of layer 5
(f) Head 4 of layer 6
(g) Head 8 of layer 7
(h) Head 8 of layer 8
(i) Head 13 of layer 9
(j) Head 12 of layer 9
(k) Head 8 of layer 11
(l) Head 5 of layer 12
Fig. 1: The decoder-only architecture model contains various attention modules with different functionalities that are not always diagonal. We selected a sentence for inference and obtained the attention maps for each head of every layer. In the figure, one head from each decoder layer is randomly selected for illustration.
@@
(c) AEAM in a repeat synthesis.
(d) AEAM in a skipping synthesis.
Fig. 2: Illustration of real attention alignments. (a), (c) and (d) are attention heatmaps from the same AEAM of the same model, with different inputs. (a) is for a text-speech pair directly picked from the test set. With the same inputs, (b) shows the reference alignment produced by the Viterbi algorithm. (c) and (d) illustrate the AEAM when some text tokens are repeated or skipped in synthesis.
In our initial experiments on the popular decoder-only TTS model VALL-E, we observed from the attention maps that certain attention heads within a specific layer exhibit diagonal patterns, indicating the alignment between speech and text. The attention maps here specifically represent the attention from speech to text, where speech tokens function as queries while text tokens act as keys. We call the attention maps of those heads Alignment-Emerged Attention Maps (AEAMs, cf. Fig. 2(a)). In the circumstances of repeating and skip generation, the alignments on AEAMs also exhibit corresponding patterns (cf. Fig. 2(c) and 2(d)). Based on this discovery, we introduce a novel inference method named Attention-Constrained Inference (ACI), which facilitates monotonic synthesis without changing the model structure and the training process. It first identifies all AEAMs among the attention maps of all layers and heads using the Attention Sweeping algorithm and then applies constraining masks (CMasks) on AEAMs to guide the monotonic generation process. The main contributions of our work are listed below:
@@ 2.2 Attention Sweeping
$\mathcal{C}_{\mathrm{E}}(\mathbf{M})=-\frac{1}{T}\sum_{t=1}^{T}\sum_{\ell=1}^{L}M_{t,\ell}\cdot\log{M_{t,\ell}}.$ (1)
@@
Fig. 3: An example of AEAM attention values and DP values, and corresponding CMasks. Here $\bm{x}$ is the input text tokens, and $\bm{y}$ is the speech tokens. (a) presents an example of the attention value pattern of a normalized AEAM when generating $y_{t+1}$, with the maximum value of each row highlighted. The values are normalized so that each row sums up to $1$. The constraining mask (CMask) for the newly generated token $y_{t}$ is now needed. (b) shows the CMask obtained via the trivial argmax strategy, which is prone to causing instability, such as the attention center being falling back. Based on attention values in (a), subfigure (c) illustrates the DP values $d$ of each state. (d) is the CMask obtained via the DP strategy, which shows stability in maintaining the monotonicity.
@@
$d_{t,\ell}=\begin{cases}d_{t-1,1}+(m_{t}-1)^{2}&\ell=1,\\ \min\{d_{t-1,\ell},d_{t-1,\ell-1}\}+(m_{t}-\ell)^{2}&l\geq 2.\end{cases}$ (2)
@@
$\mathcal{C}_{\mathrm{A}}(\mathbf{M})=\frac{1}{T}\left(\mathcal{E}(\bm{m},\bm{a},T)+\min_{c}\mathcal{E}(\bm{a}+c,\bm{b},T)\right)$ (3)
@@
$\mathcal{C}_{\mathrm{E}}\left(\mathbf{M}\right)+\mathcal{C}_{\mathrm{A}}\left(\mathbf{M}\right)<2\tau.$ (4)
@@ 2.3 Attention Constraining
VALL-E's AR synthesis often encounters issues with word skips, repeats, or pronunciation errors. We notice that when the text and synthesized speech do not perfectly match, the attention alignment in AEAMs show correspoding patterns (cf. Fig. 2(c) and 2(d)). Assuming that $y_{1}$ to $y_{t}$ have been input or synthesized and that $y_{t+1}$ is to be generated, a possible AEAM value pattern is shown in Fig. 3(a). Note that the key in the last row is the newly generated $y_{t}$, for which the alignment position needs to be calculated in this AR iteration. However, the AEAM may mistakenly align $y_{t}$ with a phoneme identical to the true one but positioned differently, leading to a non-monotonic attention alignment that misguide the subsequent generation.
@@
Fig. 4: Calculate $\mathcal{C}_{\mathrm{E}}$ and $\mathcal{C}_{\mathrm{A}}$ for each attention map of 4 models, and plot them as 2-D points. The separation line is set at $\tau=1$.
@@
Attn. Constrained AEAM(s) WER(%) $\downarrow$ Del(%) Ins(%) Sub(%)
(2-16) and (2-4) 4.14 0.65 0.65 2.84
(2-16) 4.28 0.68 0.71 2.89
all 2-nd layer attention maps 4.76 0.89 0.80 3.06
3 Experiment
@@ 4 Conclusions
Our paper introduced a training-free method, ACI, to improve the decoder-only TTS model VALL-E's synthesis robustness. Utilizing Attention Sweeping, we identify and filter AEAMs, strongly related to the alignment, then employ attention constraining to mitigate synthesis errors. Experimental results demonstrate noticeable improvements in WER without degrading naturalness. Our work demonstrates that decoder-only TTS models can unsupervisedly learn alignment-like structures and shows a way to locate and exploit them, laying the foundation for various potential works, such as streaming generation or integrate additional signals into attention maps during training.